# A method for classification of data with uncertainty using hypothesis testing

### v1

@@ 5 Discussion and conclusions
In this paper, we propose an approach to classification that incorporates uncertainty by introducing two types of hypothesis testing. In previous studies, the quantification of uncertainty required the use of resampled or validation data, as well as the tuning of multiple hyperparameters. In contrast, the proposed method determines thresholds using only training data, significantly reducing computational and time costs compared to conventional methods. Furthermore, as described in Section 2, the proposed approach establishes thresholds based on the significance level defined by hypothesis testing, providing a solid theoretical foundation. Another feature of the proposed method is its ability to leverage histograms to visually capture ambiguous regions between classes while effectively detecting uncertain data. While this study focuses on applications in the medical field, the proposed method is expected to be applicable to other domains as well.

### v2

@@ 5 Discussion and conclusions
In this paper, we propose an approach to classification that incorporates uncertainty by introducing two types of hypothesis testing. In previous studies, the quantification of uncertainty required the use of resampled or validation data, as well as the tuning of multiple hyperparameters. In contrast, the proposed method determines thresholds using only training data, significantly reducing computational and temporal costs compared to conventional methods. Furthermore, as described in Section 2, the proposed approach establishes thresholds based on the significance level defined by hypothesis testing, providing a solid theoretical foundation. Another feature of the proposed method is its ability to leverage histograms to visually capture ambiguous regions between classes while effectively detecting uncertain data.
@@
While this study focuses on applications in the medical field, the proposed method is expected to be applicable to other domains as well.